Gases are most likely to be found in a phase diagr
Answer #1

Answer 16) Correct answer is option B: High Temperature and low pressure.

SelidLiquid ree poinrt Pessure → Tripe point. Subtimation Gas

Answer 17) t1/2= 0.77h, Ao= 160 g and A= 1.59 g calculate t=?

                                   A= Ao × 2-(t/t1/2)

                                    1.59 = 160 × 2-(t/0.77)

                                ln(1.59/160) = -t/0.77 × ln 2

                                          -4.611 = -0.9t

                                              t= 5.12 h